OpenStudy (anonymous):

Bryan has a square parcel of land that is 120 yards long and 120 yards wide. The land is planted with Bermuda grass and surrounded by a fence. So that his cattle will not ruin the grass, Bryan decides to split the field into two halves and rotate his cattle on the land every 30 days. He plans to build a fence diagonally across the field. How many feet of fencing will Bryan need to split the parcel of land in half? Round your answer to the nearest yard.

OpenStudy (anonymous):

In looking at this problem it seems breaking it down into triangles would help you solve for this.

OpenStudy (anonymous):

Because you cut the square in half it is inferred that the corners of the new triangle are 45 degrees each, making the sides 45,45,90 Pythagoreums theorm states that a^2=b^2=c^2 you get 120^2+120^2=c^2

OpenStudy (anonymous):

Do that bottom math and try and try to find what c^2 is equal to and then take the square root of both and you will find "c" and check back with your answer so I can see if you followed it properly